What is the answer to 4x+y=-8 and 3x+3y=3 using elimination?
Answer by mananth(16946)

4 x + 1 y = -8 .............1

3x+3y=3
/3
1 x + 1 y = 1 .............2
Eliminate y

multiply (1)by -1
Multiply (2) by 1
-4 x -1 y = 8
1 x + 1 y = 1
Add the two equations
-3 x = 9
/ -3
x = -3
plug value of x in (1)
4 x + 1 y = -8
-12 + y = -8
y = -8 + 12
y = 4
y = 4
x= -3
y= 4
m.ananth@hotmail.ca